OpenStudy (anonymous):

An error is made in solving the equation below. What is the correct solution to the equation? 5(x – 5) = 15 5(x – 5 + 5) = 15 + 5 5x = 20 5x ÷ 5 = 20 ÷ 5 x = 4 –50 –2 8 80

OpenStudy (anonymous):

plz help guys

OpenStudy (welshfella):

the error is in the second line the correct way to start this solution would be to distribute the 5 over the x- 5

OpenStudy (welshfella):

5x - 25 = 15 can you finish this?

OpenStudy (anonymous):

so what is the solution answer?

OpenStudy (welshfella):

add 25 to both sides 5x - 25 + 25 = 15 + 25 5x = 40 x = ?

OpenStudy (welshfella):

to find x divide both sides of the '=' by 5

OpenStudy (welshfella):

was it something I said!!??

OpenStudy (anonymous):

?

OpenStudy (welshfella):

only kidding.. caan you find what x is? 5x = 40 x = ?